OpenStudy (anonymous):

Find the variance and standard deviation. Round your answers to the nearest tenth; write your answers separated by commas; example: 14.0; 5.8 {10, 12, 14, 15, 18, 20, 23}

OpenStudy (anonymous):

\[\sigma=\sqrt{\frac{\sum_{i=1}^{n}{x_{i}^2}}{n}-\mu^2}\]Just square whatever you get for that to find the variance.

OpenStudy (anonymous):

didnt understand one thing.

OpenStudy (anonymous):

You're not familiar with that notation?! Basically, what it says is: square each value, then add them all up, then divide by the number of values, then subtract the mean.

OpenStudy (anonymous):

mean squared*

OpenStudy (amistre64):

sum all the: (x-u)^2 and divide by the number in the population to get variance right?

OpenStudy (anonymous):

Yeah but the formula I put up there is quicker.

OpenStudy (amistre64):

thats why they call it a shortcut :)

OpenStudy (anonymous):

Yup :)

OpenStudy (amistre64):

my ti83 say: ox = 4.24...

OpenStudy (amistre64):

so o^2 = 17.64 or so?

OpenStudy (anonymous):

I got variance as 18.